An ampere 4a Two resistors 6 and 18 ohms are connected in parallel. Find the potential difference across the 18ohm resistor


1
Expert's answer
2019-10-07T10:53:59-0400

Let "i_1" be the current through 18 ohm resistance and

"i_2" be the current through 6 ohm resistance

For parallel connection potential defferene remain same

So,

"18i_1=6i_2\ \ \ \ \ \ \ .......(1)\\also\\i_1+i_2=4\ \ \ \ \ \ \ .......(2)"

Putting the value of "i_2" from (2) into (1)

Then,

"18i_1=24-6i_1\\24i_1=24\\or\\i_1=1\ A"

Potential difference across 18 ohm resistance

"=18i_1=18×1=18\ V"